### v1

# Efficient Multi-Agent Collaboration with Tool Use for Online Planning in Complex Table Question Answering

Abstract
Complex table question answering (TQA) aims to answer questions that require complex reasoning, such as multi-step or multi-category reasoning, over data represented in tabular form. Previous approaches demonstrated notable performance by leveraging either closed-source large language models (LLMs) or fine-tuned open-weight LLMs. However, fine-tuning LLMs requires high-quality training data, which is costly to obtain, and utilizing closed-source LLMs poses accessibility challenges and leads to reproducibility issues. In this paper, we propose Multi-Agent Collaboration with Tool use (MACT), a framework that requires neither closed-source models nor fine-tuning. In MACT, a planning agent and a coding agent that also make use of tools collaborate to answer questions. Our experiments on four TQA benchmarks show that MACT outperforms previous SoTA systems on three out of four benchmarks and that it performs comparably to the larger and more expensive closed-source model GPT-4 on two benchmarks, even when using only open-weight models without any fine-tuning. We conduct extensive analyses to prove the effectiveness of MACT’s multi-agent collaboration in TQA.
1 Introduction
The goal of table question answering (TQA) is to answer a question based on data represented in tabular form, optionally also using additional textual context. Recent studies on TQA focus more and more on complex instances, as they are ubiquitous in table data analysis (Zhu et al., 2021; Zhang et al., 2024b; Lu et al., 2023). Solving those complex instances requires performing multiple reasoning steps and/or employing different reasoning strategies (Ghosal et al., 2023). We refer to these aspects as “multi-step” and “multi-category” reasoning, respectively. An example requiring both multi-step and multi-category reasoning is shown in the upper left part of Figure 1. To answer the question, a system needs to first use factual knowledge to extract countries in Europe. Then, numerical reasoning is applied to calculate the percentage change and carry out the comparison.
Figure 1: Overview of MACT, an iterative collaboration framework for TQA that consists of five stages for each iteration as well as an efficiency optimization module.
One popular approach for addressing those complex instances in TQA is planning, where step-wise plans are generated and used to guide the reasoning process (Zhang et al., 2024c; Wang et al., 2024; Wu and Feng, 2024; Zhu et al., 2024; Zhao et al., 2024). State-of-the-art works in this direction either fine-tune open-weight large language models (LLMs) Wu and Feng (2024); Zhu et al. (2024) or prompt closed-source commercial LLMs Wang et al. (2024); Zhang et al. (2024c); Zhao et al. (2024). However, fine-tuning requires high-quality data, which is usually hard to obtain, while prompting closed-source commercial LLMs can be costly and poses challenges to reproducibility. Furthermore, these methods rely on a single LLM, hence they struggle with performing different types of reasoning. To the best of our knowledge, existing models leverage a single LLM to perform planning and reasoning, which is sub-optimal in particular if the LLM does not excel at mathematical reasoning or coding (Wu and Feng, 2024).
To address these challenges, we propose a multi-agent collaboration framework with tool use. Our framework neither depends on closed-source LLMs nor requires fine-tuning. In fact, its backbone LLMs can be exchanged flexibly. It incorporates two agents (a planning agent and a coding agent) and a set of tools (a Python interpreter, a calculator and Wikipedia search). The planning agent performs online planning, i.e., it generates a plan iteratively. This breaks down complex problems and helps address multi-step reasoning. The coding agent and the tool set assist with generating faithful intermediate results. We let these agents work in a collaborative setting, addressing the challenges of multi-category reasoning as all agents can concentrate on the reasoning types they excel in. We further introduce an efficiency optimization module which allows the framework to take shortcuts. We conduct experiments on four popular TQA benchmarks that include complex TQA instances. Our framework outperforms previous SoTA systems on three out of four benchmarks. It achieves comparable results to GPT-4 on two benchmarks even when using only open-weight models without any fine-tuning. In comparison to fine-tuned SoTA TQA systems, it demonstrates considerably better generalizability. Our analysis proves the effectiveness of our proposed collaborative setting of specialized agents. We further show that the efficiency optimization module can save up to 33% of iterations without performance degradation.
2 Related Work
We review previous work for three core aspects of MACT: planning, multi-agent collaboration and LLMs with tool use. Table 1 compares MACT with previous TQA systems.
System Online planning  No fine- tuning NL plan Multi agents Tools
Raven - - cal, SQL
Binder - - SQL/Python
Lever - - SQL
TableLlaMA - -
Dater
TAT-LLM cal
Chain of table
Reactable SQL, Python
Protrix SQL
TAPERA Python
MACT (ours) Python, cal, Wiki
Table 1: Comparing MACT with previous works. NL plan stands for using natural language to encode a plan or not. cal and Wiki refer to calculator and Wikipedia.
Planning.
We categorize previous work into three groups based on planning strategies: Heuristic coarse-grained planning consists of the same two steps retrieving and aggregation for all questions Ye et al. (2023); Zhou et al. (2024). Online global planning generates a plan in the first iteration and revises it in the next one Zhao et al. (2024). Online iterative planning conditions the generation of the next step on the executed results of previous steps Zhang et al. (2023a); Wang et al. (2024). We opt for online iterative planning for complex TQA, as it introduces more granularities (in contrast to heuristic planning) and emphasizes the dependency among steps (in contrast to online global planning). Focusing on the dependency among steps allows a system to look into a problem in depth. This aligns with the task profile of complex TQA. In contrast to previous work using iterative planning, we introduce an efficiency optimization module to minimize the costs of the framework. To learn plan generation, previous work either depends on fine-tuning (Wu and Feng, 2024; Zhu et al., 2024), or strong closed-source models, combined with in-context learning (Wang et al., 2024; Zhang et al., 2023a; Zhao et al., 2024). By contrast, MACT is suitable for both closed-source and open-weight models, without the need for fine-tuning.
Multi-agent collaboration.
In multi-agent collaboration settings, multiple AI entities collaborate towards a common goal (Talebirad and Nadiri, 2023). We use the term agent to refer to LLMs that can interact with executable tools, following Qiao et al. (2024). As far as we know, none of the previous works in TQA utilize multi-agent collaboration. The most related works explore collaboration among the same set of agents, i.e., all agents use the same backbone but are prompted differently (Liu et al., 2023; Zhao et al., 2024). The effectiveness of this approach relies heavily on the strong backbone models, which usually are closed-sourced. Our work explores multi-agent collaboration for TQA, without any constraint of model type.
Tool use.
LLMs have been shown to be ineffective in retrieving information from long tables (Zhou et al., 2024) and carrying out numerical reasoning (Imani et al., 2023). Making use of tools can ensure faithful results of these operations. The most common tools used in TQA are SQL interpreters Cheng et al. (2022), Python with Pandas dataframes Gemmell and Dalton (2023), and calculators (Zhu et al., 2024). In MACT, we use similar tools. Inspired by Shinn et al. (2023), we further add Wikipedia search as an additional tool to assist questions requiring factual knowledge.
3 Method
We propose MACT, a Multi-Agent Collaboration framework enriched with a set of Tools for TQA. Figure 1 provides an overview of the framework. It consists of four major modules: a memory $S$, a planning agent $M_{p}$, a coding agent $M_{c}$, and a tool set $T$. $M_{p}$ and $M_{c}$ are instantiated by (potentially) different LLMs. They collaborate through five core stages: action generation, action selection, tool selection/code creation, observation computation, and memory state update. The stages are executed iteratively for a maximum of $I$ iterations, where $I$ is a hyper-parameter. We control the overall efficiency of the collaboration via an efficiency optimization module. For a new TQA instance, we initialize the memory state $s_{0}=(\text{table},\text{question},\text{texts})$, i.e., with the input table, the question, and (if given) textual context. [^1] Some TQA datasets, such as TAT, provide textual context in addition to tables and questions. All parts of the memory are represented as strings. To represent the table as a string, we use pipes as column separators.
3.1 Action Generation
To format our plans, we follow ReAct (Yao et al., 2022) that consists of the generation of thoughts, actions and observations. Our framework only requires actions and observations, but following Yao et al. (2022), who demonstrate performance gains from generating thoughts with actions, we adopt their prompting method. Thus, at each iteration $i\leq I$, we prompt $M_{p}$ to generate a thought $z_{i}$, an action $a_{i}$ and an observation $\hat{o}_{i}$: $(z_{i},a_{i},\hat{o}_{i})\sim M_{p}(z_{i},a_{i},\hat{o}_{i}|s_{i-1},\phi_{p},\tau_{p})$, where $s_{i-1}$ is the previous memory state, and $\phi_{p}$ and $\tau_{p}$ are the prompt (provided in A.6) and temperature of the LLM used for $M_{p}$, respectively. Note that $\hat{o}_{i}$ is not the final observation. Instead, it is later used during execution as a particular form of our proposed collaboration between the planning and coding agent (see 3.5). We sample from $M_{p}$ $k$ times, resulting in $k$ actions $\{a_{i}^{n}\}_{n\leq k}=\{a_{i}^{1},a_{i}^{2},...,a_{i}^{k}\}$ and their corresponding estimated observations $\{\hat{o}_{i}^{n}\}_{n\leq k}=\{\hat{o}_{i}^{1},\hat{o}_{i}^{2},...,\hat{o}_{i}^{k}\}$ in iteration $i$. Following Yao et al. (2022), we define an action $a_{i}$ with two parts: an intent and an instruction, e.g., Retrieval [Retrieve the export number for France and Germany].” The intent encodes the purpose of an action, e.g., Retrieval denotes retrieving information from the input table. The instruction (marked with brackets) provides detailed specifications of the intent. Table 2 shows the six types of intents we define for our framework and examples for corresponding instructions.
Intent Instruction: Format and Example Tool Selection Code Generation Tool Use/Execution
Retrieval textual description of what to retrieve, e.g., “sell numbers of 2019” $t=Python$ $M_{c}$ Python interpreter is run on generated code
Calculation formula or textual description of what to calculate, e.g., “(135-114)/135” $t=T_{cal}$ if formula, else $t=T_{python}$ $M_{c}$ if not formula Calculator is executed or Python interpreter is run on code
Search entity name, e.g., “Tesla” $t=T_{srh}$ no Wikipedia API is called on entity
Read textual description of required information from input texts, e.g., “when was the target method adopted” $t=Null$ no $M_{p}$ is prompted to extract information from provided textual context
Finish final answer $y$ $t=Null$ no final answer is output and execution stops
Ask textual description of required information from $M_{p}$, e.g., “hours of a day” $t=Null$ no $M_{p}$ is prompted for the information need
Table 2: Overview of intents and instructions of actions and how they are executed within our framework.
The intents Retrieval and Calculation are commonly seen in previous works (Gemmell and Dalton, 2023; Zhu et al., 2024). We use Retrieval for any operations extracting information from a table, including direct querying, filtering, and grouping. Instructions that require calculation, counting or comparison are captured by Calculation. To fulfill the possible need for external (factual) knowledge that is not present in the table or textual context, we add the intent Search. Read covers the need for contextual reasoning in table-text QA. It refers to instructions involving retrieving information from the texts provided as part of TQA instances. The intent Finish stops $M_{p}$ from generating more actions and ends the iterative execution of our framework, providing the final answer in the corresponding instruction. Lastly, we use an intent called Ask to retrieve an answer based on the internal knowledge from the planning agent.
3.2 Action Selection
From the set of $k$ actions generated by $M_{p}$ for iteration $i$, we use a function $f_{s}$ to select the most promising action $a^{*}_{i}=f_{s}(s_{i-1},\{a_{i}^{n}\}_{n\leq k})$. We use self-consistency (SC) (Wang et al., 2022) as the selection function, which outputs the most frequent action from the set of sampled actions. In the case of ties, we choose the most frequent action that was sampled first. We provide a comparison with other selection functions in A.4.
3.3 Tool Selection and Use
The tool needed for executing action $a^{*}_{i}$ depends on the intent of the action (see columns “Tool Selection” and “Tool Use/Execution” of Table 2). To address the intents Search, Calculation and Retrieval, we introduce a set of tools $T=\{T_{srh},T_{cal}\,T_{python}\}$. $T_{srh}$ is an API function for Wikipedia search (Shinn et al., 2023) from Langchain. [^2] https://rubydoc.info/gems/langchainrb/Langchain/Tool/Wikipedia The API takes a target entity specified in the instruction and returns the first paragraph of the corresponding Wikipedia entry. $T_{cal}$ is a calculator, powered by a Python interpreter. It takes a formula generated by $M_{p}$ and outputs the answer. Note that the instruction of Calculation can also be a textual description, such as “Compute the average number of medals for each country in the table.” To better address these instructions, we introduce a coding agent $M_{c}$ with a Python interpreter, denoted as $T_{python}$ (see 3.4). The intent Retrieval is also addressed by $M_{c}$ and $T_{python}$, i.e., $M_{c}$ generates Python code based on a given instruction to retrieve target cells in the tables and the Python interpreter returns the executed results. Lastly, For Read, Ask and Finish, no tool is used, denoted as $t=Null$ in Table 2. The answers to the intents Read and Ask are queried via $M_{p}$. For Read, $M_{p}$ reads the answer from a given textual context. For Ask, it responds based on its internal knowledge. No execution is performed for Finish, as the intent ends MACT with a final answer in its instruction.
3.4 Code Generation and Execution
To address textual instructions for Calculation actions as well as Retrieval actions, we integrate a coding agent $M_{c}$, which is an LLM and translates the instructions of $a^{*}_{i}$ into Python code snippets $c_{i}\sim M_{c}(c_{i}|a^{*}_{i},s_{i-1},\phi_{c},\tau_{c})$. The hyper-parameter $\tau_{c}$ controls the temperature of the coding agent, and $\phi_{c}$ is a static, pre-defined prompt (see A.6). We sample $k$ times from $M_{c}$ to increase the robustness of the system against generated syntax errors, resulting in a set of code snippets $C=\{c_{i}^{n}\}_{n\leq k}$. A Python interpreter is run on each $c_{i}$, creating a set of executed solutions $\hat{C}=\{\hat{c}_{i}^{n}\}_{n\leq k}$
3.5 Observation Computation
The computation of the final observation $o_{i}$ depends on the selected tool $t_{i}$: if $t_{i}\in\{T_{cal},T_{srh}\}$, the corresponding tool returns a deterministic result. If $t_{i}$ is $T_{python}$, we select the most frequent element from the combined set $\{\hat{o}_{i}^{n}\}_{n\leq k}\cup\hat{C}$ as final observation, where $\{\hat{o}_{i}^{n}\}_{n\leq k}$ is the estimated observations sampled from $M_{p}$. This strategy features two levels of collaboration: from an ensemble perspective, both $M_{p}$ and $M_{c}$ contribute to obtain $o_{i}$; from a pipeline perspective, $M_{c}$ makes use of the outputs (actions) of $M_{p}$. If neither $T$ nor $M_{c}$ are needed to execute the action, the final observation is the most frequent element in $\{\hat{o}_{i}^{n}\}_{n\leq k}$.
3.6 Memory State Update and Iteration
After obtaining $o_{i}$, we update the memory state with the selected action and the observation of iteration $i$: $s_{i}=s_{i-1}+[a^{*}_{i},o_{i}]$. The framework then continues with the next iteration $i+1$. Note that adding the observation to the memory state also allows $M_{p}$ to build on top of results from $M_{c}$ in iteration $i+1$. If $i>I$, the execution stops with a predicted answer directly from $M_{p}$. In 98% of the cases, a final answer is given before $i>I$.
3.7 Efficiency Optimization
The iterative collaboration approach of $M_{p}$ and $M_{c}$ is highly effective in practice as demonstrated in our experiments. However, questions that do not require multi-step or multi-category reasoning can also be answered directly by $M_{p}$. For those instances, we propose an efficiency optimization component that serves as a shortcut for directly outputting an answer in the first iteration.Whether the answer is output directly depends on the confidence of $M_{p}$, which we approximate by the degree of self-consistency of its estimated predictions $Y=\{y^{1},...,y^{k}\}$. $Y$ is obtained by accessing the whole reasoning trace (consisting of $j$ actions and estimated observations) that $M_{p}$ generates for a given $s_{0}$ until the intent Finish is output: i.e., $(a^{k}_{1},\hat{o}^{k}_{1},...a^{k}_{j},\hat{o}^{k}_{j})\sim M_{p}(a^{k}_{1},\hat{o}^{k}_{1},...a^{k}_{j},\hat{o}^{k}_{j}|s_{0},\phi_{p},\tau_{p})$. The output $y^{k}$ is the instruction of the action $a^{k}_{j}$ with intent Finish. To control the trade-off between performance and computation time, we introduce a hyper-parameter $\alpha\in[0..1]$. If the degree of self-consistency, i.e., the number of occurrences of the most frequent prediction in $Y$ is larger than $\alpha*k$ (high degree of SC), $M_{p}$ outputs this most frequent answer. Otherwise, the collaborative framework as described before is adopted. In short, the smaller $\alpha$, the more often the system is allowed to use the shortcut, and the larger $\alpha$, the more confident $M_{p}$ needs to be in order to take the shortcut.
4 Experiments
We assess the performance of MACT on four TQA benchmarks in comparison to SoTA TQA systems.
Datasets.
We choose four TQA datasets that cover different reasoning complexities and domains (See A.1 for more details). WTQ Pasupat and Liang (2015) is the easiest dataset as it neither requires multi-step nor multi-category reasoning. However, it is a widely used benchmark for TQA in the general domain and enables a fair comparison with recent TQA systems. TAT (Zhu et al., 2021) includes hybrid tabular and textual data. Most questions require numerical reasoning. CRT (Zhang et al., 2023b) uses Wikipedia tables and involves complex reasoning. SCITAB (Lu et al., 2023) contains claims requiring compositional reasoning for verification. We follow the original work to convert it from the setting of fact verification to TQA.
TQA systems for comparison.
We categorize the recent TQA systems into two groups indicating if they require LLM fine-tuning or not. As systems requiring fine-tuning, we include OmniTab (Jiang et al., 2022), TableLlama Zhang et al. (2024a), ProTrix (Wu and Feng, 2024), TAT-LLM (Zhu et al., 2024) and TableLLM Zhang et al. (2024b). Except for OmniTab, which is backboned by BART (Lewis et al., 2020), all others build on top of LLaMA 7b (Touvron et al., 2023). As TQA systems that do not require fine-tuning, we include Dater (Ye et al., 2023), Binder (Cheng et al., 2023), Chain-of-Table (Wang et al., 2024), ReAcTable (Zhang et al., 2024c), TabSQLify (Nahid and Rafiei, 2024), Plan-then-Reason (Wu and Feng, 2024), Mix-SC (Liu et al., 2023) and ARC Zhang et al. (2023b). They all use GPT-3.5-turbo.
Experimental settings.
In MACT, the choice of the planning and coding agent is flexible as no fine-tuning is involved. We experiment with the best open-weight LLMs available at the time of writing: Qwen 72B (Yang et al., 2024) (planning agent) and CodeLLaMA-34B (coding agent). They are run on int4 and full precision, respectively. We further use GPT-3.5-turbo as both the planning and coding agents when comparing our method with other TQA systems that use this model. $\tau_{p}$ and $\tau_{c}$ are set to 0.6 for non-repetitive action generation. We set the action and code generation size $k$ to 5, following Liu et al. (2023). The maximum number of iteration $I$ is set to 7, based on empirical results on the development sets. For the efficiency component, we set $\alpha$ to 1 to ensure high confidence of the model. We explore the effects of different values of $\alpha$ in Section 6. Two NVIDIA A100 GPUs are used for running MACT.
5 Results
We first evaluate the performance of MACT in direct comparison to recent TQA systems using closed-source LLMs. Second, we examine how our method performs compared to fine-tuned open-weight LLMs. Following prior work, we use exact match (EM) as the evaluation measure. Lastly, we discuss the efficiency of MACT.
MACT outperforms TQA models on three out of four datasets when using GPT-3.5 as the backbone.
The upper part of Table 3 shows MACT using GPT-3.5 as the planning and coding agent in comparison to SoTA TQA systems using GPT-3.5 as the backbone LLM. MACT (GPT-3.5) surpasses the examined TQA models, except for Mix-SC on WTQ. This indicates the effectiveness of our multi-agent strategy compared to single-agent TQA models. We suspect that the performance gap between our approach and Mix-SC comes from data-specific table-cleaning and answer format controls in Mix-SC. In contrast, MACT does not include any dataset-specific pre- or postprocessing steps to keep it generally applicable to any dataset.
WTQ TAT CRT SCT
closed-source LLM backbones
GPT-3.5 45.8 39.7 39.3 48.9
Dater 52.8* 22.1 46.8 47.1
Binder 56.7* 0.9 1.24 29.1
Chain-of-Table 59.9* 20.5 33.9 27.6
ReAcTable 52.4* 9.26 29.8 32.1
TabSQLify 64.7* 13.7 42.0 50.9
Plan-then-Reason 65.2* 41.2 44.9 52.5
Mix-SC 73.6* 54.3 48.6 49.3
ARC - - 56.3* -
MACT 70.4 64.5 57.4 55.8
open-weight LLM backbones
Qwen (Qw-72b) 60.6 53.6 55.9 55.0
CodeLLaMA (CL-34b) 55.0 29.5 49.7 9.5
SC(Qw-72b+CL-34b) 69.0 56.7 61.4 54.4
MACT (Qw-72b+Qw-72b) 68.6 66.3 59.8 57.3
MACT (CL-34b+CL-34b) 55.2 54.1 43.5 45.0
MACT (Qw-72b+CL-34b) 72.6 66.2 64.4 59.8
GPT-4 72.9 80.8 58.7 63.2
Humans (Crowdsourcing) - 84.1 - 84.7
Table 3: Results of models without fine-tuning. SCT refers to SCITAB. The models are grouped by the LLM they use as their backbone (top: GPT-3.5; bottom: open-weight LLMs as indicated in parentheses). Performances marked with * are taken from the original paper. Performances marked with †are taken from Wu and Feng (2024), Zhu et al. (2024), Zhang et al. (2023b) and Lu et al. (2023) for each dataset.
MACT outperforms out-of-the-box open-weight LLMs across datasets, demonstrating the effectiveness of specialized agents.
The middle part of Table 3 provides the results of MACT using specialized agents (MACT (Qw+CL): Qwen as planning agent, CodeLLaMA as coding agent). As baselines, we compare a setting without a specialized coding agent (MACT (Qw+Qw)) and a setting without a general planning agent (MACT(CL+CL)). As further baselines, we use the two LLMs on their own as well in combination, with Chain-of-thought prompting (Wei et al., 2022). For combination, the single models are prompted five times and combined using SC, as in Liu et al. (2023). This is a direct multi-agent baseline without collaboration and tool use. Both SC(Qw+CL) and MACT (Qw+CL) achieve higher EM scores than individually prompting Qwen and CodeLLaMA, demonstrating the positive effect of using multiple agents for planning and coding. Importantly, MACT (Qw+CL) outperforms SC(Qw+CL) by approximately 6 EM points on average across all datasets, highlighting the superiority of our collaboration technique over simply taking the most frequent predictions from two independent agents. We also find that having an expert coding agent for code generation (MACT (Qw+Qw) vs. MACT (Qw+CL)) improves performance considerably.
c WTQ TAT CRT SCT
OmniTab (BART 406m) c 62.3* 17.1 20.6 29.1
TableLlama (LlaMA 7b) c 29.9 17.4 26.9 38.6
Protrix (LlaMA 7b) c 48.9* 26.8 40.2 42.4
TAT-LLM (LlaMA 7b) c - 69.6* - -
MACT(LlaMA-7b+CS-7b) c 38.1 28.3 40.0 41.1
MACT(Qw-7b+CS-7b) c 58.4 61.9 46.4 45.9
Table 4: Results of MACT using different LLM agents in comparison to fine-tuned TQA models. Performances marked with * refer to the in-domain setting (where fine-tuning took place). SCT refers to SCITAB. CS refers to the deepseek-coder model.
MACT with open-source models delivers comparable performances as closed-source systems.
As shown in Table 3, by comparing MACT (Qw+CL) with TQA systems that rely on closed-source LLMs (in the upper part of Table 3), we find that our model outperforms the examined TQA systems for three out of four datasets. This is an important finding, as TQA systems based on commercial closed-source LLMs are costly to operate and often lack reproducibility. In contrast, our multi-agent TQA system is more cost-efficient and straightforward to replicate, while delivering superior performance compared to closed-source systems. In addition, we show two upper-bounds in the bottom part of Table 3: The performance of human annotators and of GPT-4. We observe that MACT achieves comparable performance as GPT-4 on two datasets, without the need for fine-tuning.
MACT generalizes better across datasets than fine-tuned TQA systems.
Table 4 compares our framework with prior fine-tuned TQA models. We present results for MACT with different planning and coding agents: Our standard setting (LlaMA-7b+CS-7b) as well as with a stronger planner (Qw-7b+CS-7b). For the models from related work, we mark the dataset on which they have been fine-tuned with * signs. In general, their performance on the dataset used for fine-tuning is rather high while they suffer from a considerable drop in EM when tested on other datasets. This observation is in line with the findings by (Zhang et al., 2024a; Huang et al., 2024). In contrast, MACT does not use fine-tuned models and can, thus, be applied to any dataset with a good generalization performance. MACT demonstrates comparable results as Protrix when using LlaMA-7b as the planning agent, though it has not been fine-tuned. As expected, using a better planning agent leads to better results. This also shows the robustness of MACT in terms of backbone models.
MACT is moderately efficient.
Table 5 compares MACT with other approaches in terms of the total number of prompting (generation) for each instance. For Binder and Dater, SC is performed a fixed number of times regardless of problem complexity. This results in a high number of prompting per instance, thus being inefficient. In contrast, MACT provides flexibility in generation, as the number of iterations depends on the problem’s complexity. For instance, most questions can be solved within three steps for WTQ (see our analysis in A.3), this results in a total of 25 times prompting3332 steps involve action and execution generation, with each five times, plus last step five times of action generation: 2*(5+5)+5. at most for each instance. If we incorporate the efficiency optimization module, which potentially saves up to one-third of the iterations (see Section 6), the total number of generations per instance is even lower (approximately 15), making MACT comparable to other approaches in terms of efficiency. The iterative nature of MACT can lead to a higher upper-bound of generation. However, it also allows for more complex problems, making the approach more tailored to real-life needs.
c Prompting number per instance
Binder c 50
Dater c 100
Chain-of-Table c 1-25
ReAcTable c 15-125
Mix-SC c 10-30
MACT c 5-65
Table 5: Prompting number of different approaches. We show lower and upper-bound if not deterministic.
6 Analysis
We conduct various analyses of our framework to back up our claims and contributions. Unless mentioned otherwise, all analyses are performed using Qwen-72b as $M_{p}$, CodeLlama-34b as $M_{c}$, the number of action generation $k=5$, and selection model $f_{s}=\texttt{SC}$. To explicitly analyze the effects of multi-agent collaboration with tool use, we do not use efficiency optimization, which means all instances undergo the iterative collaboration between $M_{p}$ and $M_{c}$ with tool use. Further analysis to support our choices of the sampling size $k$ and the maximum number of iterations $I$ are in A.2 and A.3.
WTQ TAT CRT SCITAB
MACT (Qw-72b) 72.6 66.2 64.4 59.8
     w/o $T_{srh}$ 72.0 66.2 64.6 59.6
     w/o $T_{srh}+T_{cal}$ 71.3 62.8 63.9 58.2
     w/o $T$+$M_{c}$ 67.1 61.2 60.4 57.9
Table 6: Ablation study. $T_{srh}$ and $T_{cal}$ refer to the Wikipedia search tool and the calculator tool. $T$ includes the above two tools and a Python interpreter. $M_{c}$ is the coding agent.
Effect of Multi-Agent-Collaboration with Tool Use.
We explore the effectiveness of specialized agents and tool use in MACT by conducting an ablation study with three scenarios: ablating only $T_{srh}$ (Wikipedia search API), ablating $T_{srh}$ and $T_{cal}$ (calculator), and further ablating the coding agent $M_{c}$ with a Python interpreter. In cases where $M_{c}$ or/and tools are ablated, the most frequent estimated observations from $M_{p}$ are used as the final observations. Our results in Table 6 show that both the tools and the coding agent contribute to the performance of the framework. Nevertheless, they contribute differently to the final performance. For instance, ablating the search tool barely influences the results whereas there are large performance drops when further ablating the coding agent and the Python interpreter. This might be due to the frequency of the usage of tools and the coding agent. We find that the search tool is barely used whereas the coding agent is called in almost every query. Detailed analysis can be found in A.5. In addition, since Wikipedia is a common pre-trained corpus for LLMs, most information might have already been encoded. Nevertheless, the search tool can still be helpful given LLMs are known to suffer from hallucinations and the knowledge encoded might not be updated in time. We further observe that the ablation affects WTQ and TAT more than CRT and SCITAB. This might be attributed to dataset features: CRT contains many yes-no questions and SCITAB has been converted from a ternary classification dataset. Thus, chances for guessing the correct final answers are higher than in datasets with a more diverse answer distribution, such as WTQ and TAT. By evaluating our framework on instances from CRT that have answers other than yes/no, we find a performance drop of 8.23 when ablating both tools and the coding agent.
Effect of Efficiency Optimization.
Figure 2: EM (bar chart) and iteration ratio (line chart) against different $\alpha$. The iteration ratio is calculated by dividing the number of iterations when using efficiency optimization with a specific $\alpha$ by the number of iterations when not using it.
To investigate the trade-off between efficiency and accuracy, we plot the model performance against $\alpha\in\{0.2,0.4,0.6,0.8,1\}$. We also plot the ratio of the total number of iterations taken to terminate from each tested $\alpha$ to the total number of iterations taken when not using the optimization, i.e., letting the planning agent decide via the Finish action when to stop the execution. Figure 2 shows that the best performance is reached for $\alpha=1$, i.e., when requiring all estimated results to agree with each other in order to stop the iteration. For SCITAB, for instance, we save approximately 40% of the iterations when setting $\alpha$ to 1 without losing performance compared to not using the optimization component (59.8 vs. 59.7) On average, adding the efficiency optimization module saves up to 33% of iterations. This shows the effectiveness of the optimization in general and that users can individually tune the desired trade-off of performance and computation time.
Error Analysis
We randomly sample 50 instances that MACT fails per dataset and conduct an error analysis. About half of the errors come from the invalid or wrong code generated by the coding agent $M_{c}$. Either $M_{c}$ fails to understand instructions, thus generating wrong code, or due to the complex table data type, the code execution is not successful. This indicates the importance of table pre-processing. The second error type can be attributed to evaluation. We find that about one-third of failures come from strict evaluation metrics (Exact Match Accuracy). This influences the performance of MACT on the TAT dataset the most, as it features long text strings as answers. The evaluation challenge has been discussed in many previous works Wu and Feng (2024); Li et al. (2024). To estimate the upper-bound of our method, we use GPT-4 for evaluation and achieve an accuracy of 87.8% on the TAT dataset (66.2% with EM). The remaining error cases can be largely attributed to the failure of the planning agent. This means the planning agent does not decompose questions correctly. For instance, one question asks the score range (min-max) for the top 10 finishers. Apart from retrieving the min and max scores of the top 10 finishers, the planner continues to generate the action: Calculate[Calculate the range of the scores in the observation 1.]. This results in the wrong final prediction of the difference between the min and max scores.
7 Conclusions
We propose multi-agent collaboration with tool use for table question answering. Unlike previous work, our MACT framework neither requires fine-tuning nor does it depend on closed-source models. In our experiments, our framework demonstrates generalizability across different benchmark datasets and outperforms a number of state-of-the-art approaches, including closed-source commercial models and fine-tuned models. To boost efficiency, we introduce an efficiency optimization module that saves up to 33% of the iterations in our analysis. In our experiments and analyses, we show that multi-agent collaboration with tools is an effective approach for table question answering.
8 Limitation
MACT is evaluated mainly with single table settings due to the scarcity of datasets featuring multi-table complex reasoning. Though the framework can be extended easily to deal with multiple tables by concatenating them in the inputs, it is still not clear how effective our approach will be in a multi-table setting. Secondly, we only study TQA in the context of English, while there exist many multi-lingual TQA benchmarks and challenges.
Appendix A Appendix
A.1 Datasets
Table A.1 shows their statistics and characteristics. WTQ Pasupat and Liang (2015), TAT (Zhu et al., 2021), CRT Zhang et al. (2023b) and SCITAB (Lu et al., 2023) are publicly available under the licenses of CC-BY-SA-4.05, MIT and MIT, respectively. These licenses all permit us to compose, modify, publish, and dis- tribute additional annotations upon the original dataset.
Dataset #Test M-step M-category Domain
WTQ 4,344 General
TAT 1,663 Financial
CRT 728 General
SCITAB 1,162 Scientific
Table 7: We use four datasets that vary in reasoning complexity (M-step: multi-step, M-category: multi-category reasoning) and domain. #Test refers to the number of test instances.
A.2 Effect of Sampling Size
Figure 4 shows the effect of the number of generated actions $k$ on the results. Generally, we find that a larger $k$ results in better performance. However, the performance gain is small when increasing $k$ from 5 to 10. We even observe a slight performance drop for SCITAB when increasing $k$ from 5 to 10. Based on these observations, we argue $k=5$ is a good choice for the number of generated action in MACT.
A.3 Analysis of Iteration Number Distribution
We analyze the distribution of numbers of iterations for each dataset in Figure 5. Most of instances can be solved within seven iterations. Dataset-wise, CRT and SCITAB seem to require more iterations than WTQ and TAT, indicating their difficulties in terms of multi-step reasoning.
A.4 Choice of Selection Model.
In MACT, we use SC as the action selection model (see Section Action Selection). We now provide results for alternative selection models that have been introduced by prior work. In particular, we compare to LLM-based selection (Yao et al., 2023), log probability (Zhang et al., 2024c), roll-out (Hao et al., 2023) and a combination of all strategies (Hao et al., 2023). In the LLM strategy, an LLM evaluator is utilized to select the best action. For consistency with our other results, we use Qwen-72b as the evaluator. We use the same prompts as the original work Yao et al. (2023). Log probability (LOG_P) has been widely used to assist sub-path selection (Zhang et al., 2024c; Hao et al., 2023). However, it can only be used for open-weight LLMs, as it requires access to log probabilities. ROLL_OUT estimates future answers by rolling out the current reasoning path and selects the action that leads to the most frequent future answer. For the COMBINED method, we use majority voting among all individual selection models.
WTQ TAT CRT SCITAB
sc 72.6 (2) 66.2 (2) 64.4 (1) 59.8 (1)
llm 70.7 (4) 66.2 (2) 58.4 (5) 56.8 (4)
log_p 70.1 (5) 64.9 (5) 61.4 (3) 57.2 (3)
roll_out 71.9 (3) 65.9 (4) 60.2 (4) 55.5 (5)
combined 72.7 (1) 66.6 (1) 62.6 (2) 57.4 (2)
Table 8: Results using different selection models. We put the relative ranking of the models per dataset in parentheses.
The results in Table 8 show that for WTQ and TAT, SC and COMBINED lead to the best performance. For CRT and SCITAB, SC outperforms COMBINED, caused by the comparably poorer performance of LLM, LOG_P and ROLL_OUT on these datasets. SC is more efficient than COMBINED as the latter requires running all selection models, including the computationally expensive LLM. Overall, this analysis confirms that SC as selection model is a good choice.
A.5 Analysis of Intent Distribution
We report the distribution of the intents of the selected actions for each dataset in Figure 3. We observe that Retrieve and Calculate are the most frequent intents, along with Finish. This indicates that our proposed second agent, the coding agent $M_{c}$ is used frequently. Different datasets also require different intents. In particular, the framework needs to use the intent Read for solving instances in TAT, where textual descriptions are given while this is not the case for the other datasets. We notice that the Search intent is used only few times across datasets. This might be because most instances were designed to be solved using the given table and text information. However, when looking into the individual cases where Search is used, we still find it useful. For instance, one question from WTQ asks about the number of athletes from American but no information about nationality is given in the table. In this case, Search assists with answering the question by adding the nationality for each athlete from Wikipedia.
Figure 3: The distribution of action intents for different datasets.
A.6 MACT Prompts
We provide the prompts used for the planning agent for the examined datasets (WTQ, TAT, CRT and SCITAB) in Figures 6, 7, 8, and 9. Figure 10 and Figure 11 show the prompts used for the coding agent for the action intents Retrieval and Calculation, respectively.
Figure 4: EM against different action generation size.
Figure 5: The distribution of number of iterations for each dataset.
Figure 6: MACT: Planning agent prompt for WTQ.
Figure 7: MACT: Planning agent prompt for TAT.
Figure 8: MACT: Planning agent prompt for CRT.
Figure 9: MACT: Planning agent prompt for SCITAB.
Figure 10: MACT: Coding agent prompt for retrieval.
Figure 11: MACT: Coding agent prompt for calculation.

### v2

@@
 Complex table question answering (TQA) aims to answer questions that require complex reasoning, such as multi-step or multi-category reasoning, over data represented in tabular form.
-Previous approaches demonstrated notable performance by leveraging either closed-source large language models (LLMs) or fine-tuned open-weight LLMs.
-However, fine-tuning LLMs requires high-quality training data, which is costly to obtain, and utilizing closed-source LLMs poses accessibility challenges and leads to reproducibility issues.
-In this paper, we propose Multi-Agent Collaboration with Tool use (MACT), a framework that requires neither closed-source models nor fine-tuning.
-In MACT, a planning agent and a coding agent that also make use of tools collaborate to answer questions.
-Our experiments on four TQA benchmarks show that MACT outperforms previous SoTA systems on three out of four benchmarks and that it performs comparably to the larger and more expensive closed-source model GPT-4 on two benchmarks, even when using only open-weight models without any fine-tuning.
-We conduct extensive analyses to prove the effectiveness of MACT’s multi-agent collaboration in TQA.
+Previous approaches demonstrate notable performance by leveraging either closed-source large language models (LLMs) or fine-tuned open-weight LLMs.
+However, fine-tuning LLMs requires high-quality training data, which is costly to obtain.
+The use of closed-source LLMs poses accessibility challenges and leads to reproducibility issues.
+In this paper, we propose Multi-Agent Collaboration with Tool use (MACT), a framework that requires neither fine-tuning nor closed-source models.
+In MACT, a planning agent and a coding agent that also make use of tools collaborate for TQA.
+MACT outperforms previous SoTA systems on three out of four benchmarks and performs comparably to the larger and more expensive closed-source model GPT-4 on two benchmarks, even when using only open-weight models without any fine-tuning.
+Our extensive analyses prove the effectiveness of MACT’s multi-agent collaboration in TQA.
+We release our code publicly.
+[^1] https://github.com/boschresearch/MACT
 1 Introduction
@@
 Solving those complex instances requires performing multiple reasoning steps and/or employing different reasoning strategies (Ghosal et al., 2023).
-We refer to these aspects as “multi-step” and “multi-category” reasoning, respectively.
-An example requiring both multi-step and multi-category reasoning is shown in the upper left part of Figure 1.
-To answer the question, a system needs to first use factual knowledge to extract countries in Europe.
-Then, numerical reasoning is applied to calculate the percentage change and carry out the comparison.
+We refer to these aspects as multi-step and multi-category reasoning, respectively.
+An example requiring both types of reasoning is shown in the upper left part of Figure 1.
+To answer the question about the percentage change, a system first needs to use factual knowledge to extract countries in Europe.
+Then, numerical reasoning is applied to calculate the percentage change and to carry out the comparison.
 Figure 1: Overview of MACT, an iterative collaboration framework for TQA that consists of five stages for each iteration as well as an efficiency optimization module.
@@
 (2024).
-However, fine-tuning requires high-quality data, which is usually hard to obtain, while prompting closed-source commercial LLMs can be costly and poses challenges to reproducibility.
-Furthermore, these methods rely on a single LLM, hence they struggle with performing different types of reasoning.
-To the best of our knowledge, existing models leverage a single LLM to perform planning and reasoning, which is sub-optimal in particular if the LLM does not excel at mathematical reasoning or coding (Wu and Feng, 2024).
-To address these challenges, we propose a multi-agent collaboration framework with tool use.
-Our framework neither depends on closed-source LLMs nor requires fine-tuning.
+However, fine-tuning requires high-quality data, which is usually expensive to obtain (Zhu et al., 2021).
+Prompting closed-source commercial LLMs can also be costly and poses challenges to reproducibility.
+To the best of our knowledge, existing methods leverage a single LLM to perform planning and reasoning, which is sub-optimal in particular if the LLM does not excel at mathematical reasoning or coding (Wu and Feng, 2024).
+These models struggle with answering questions requiring complex reasoning.
+To address these challenges, we propose MACT, a multi-agent collaboration framework with tool use, which neither depends on closed-source LLMs nor requires fine-tuning.
 In fact, its backbone LLMs can be exchanged flexibly.
@@
 The planning agent performs online planning, i.e., it generates a plan iteratively.
-This breaks down complex problems and helps address multi-step reasoning.
+This breaks down complex problems and helps to address multi-step reasoning.
 The coding agent and the tool set assist with generating faithful intermediate results.
-We let these agents work in a collaborative setting, addressing the challenges of multi-category reasoning as all agents can concentrate on the reasoning types they excel in.
-We further introduce an efficiency optimization module which allows the framework to take shortcuts.
+The agents work in a collaborative setting, addressing the challenges of multi-category reasoning as all agents can concentrate on the reasoning types they excel in.
+An efficiency optimization module which allows the framework to take informed shortcuts.
 We conduct experiments on four popular TQA benchmarks that include complex TQA instances.
@@
 It achieves comparable results to GPT-4 on two benchmarks even when using only open-weight models without any fine-tuning.
-In comparison to fine-tuned SoTA TQA systems, it demonstrates considerably better generalizability.
+In comparison to fine-tuned SoTA TQA systems, it demonstrates considerably better generalizability across datasets.
 Our analysis proves the effectiveness of our proposed collaborative setting of specialized agents.
-We further show that the efficiency optimization module can save up to 33% of iterations without performance degradation.
+We find that the efficiency optimization module can save up to 33% of iterations without performance degradation.
 2 Related Work
@@
 MACT (ours) Python, cal, Wiki
-Table 1: Comparing MACT with previous works. NL plan stands for using natural language to encode a plan or not. cal and Wiki refer to calculator and Wikipedia.
+Table 1: Comparing MACT with previous works. NL plan stands for using natural language to encode a plan, cal=calculator, and Wiki=Wikipedia.
 Planning.
-We categorize previous work into three groups based on planning strategies: Heuristic coarse-grained planning consists of the same two steps retrieving and aggregation for all questions Ye et al.
+We categorize previous work into three groups based on planning strategies: Heuristic coarse-grained planning consists of two pre-defined steps of retrieving and aggregating Ye et al.
 (2023); Zhou et al.
@@
 (2024).
-We opt for online iterative planning for complex TQA, as it introduces more granularities (in contrast to heuristic planning) and emphasizes the dependency among steps (in contrast to online global planning).
-Focusing on the dependency among steps allows a system to look into a problem in depth.
-This aligns with the task profile of complex TQA.
+We opt for online iterative planning for complex TQA, as it provides for more fine-grained steps during problem solving (in contrast to heuristic planning) and emphasizes the dependency among steps (in contrast to online global planning), which is crucial in complex TQA.
 In contrast to previous work using iterative planning, we introduce an efficiency optimization module to minimize the costs of the framework.
-To learn plan generation, previous work either depends on fine-tuning (Wu and Feng, 2024; Zhu et al., 2024), or strong closed-source models, combined with in-context learning (Wang et al., 2024; Zhang et al., 2023a; Zhao et al., 2024).
-By contrast, MACT is suitable for both closed-source and open-weight models, without the need for fine-tuning.
+To learn how to generate plans, previous work either depends on fine-tuning (Wu and Feng, 2024; Zhu et al., 2024), or strong closed-source models, combined with in-context learning (Wang et al., 2024; Zhang et al., 2023a; Zhao et al., 2024).
+By contrast, MACT generates effective plans using either closed-source or open-weight models, without the need for fine-tuning.
 Multi-agent collaboration.
 In multi-agent collaboration settings, multiple AI entities collaborate towards a common goal (Talebirad and Nadiri, 2023).
-We use the term agent to refer to LLMs that can interact with executable tools, following Qiao et al.
+We use the term agent to refer to LLMs that interact with executable tools, following Qiao et al.
 (2024).
 As far as we know, none of the previous works in TQA utilize multi-agent collaboration.
-The most related works explore collaboration among the same set of agents, i.e., all agents use the same backbone but are prompted differently (Liu et al., 2023; Zhao et al., 2024).
-The effectiveness of this approach relies heavily on the strong backbone models, which usually are closed-sourced.
+The approaches most closely related to ours explore collaboration among homogenous agents, i.e., all agents use the same backbone but are prompted differently (Liu et al., 2023; Zhao et al., 2024).
+The effectiveness of this approach relies heavily on strong (closed-source) backbone models.
 Our work explores multi-agent collaboration for TQA, without any constraint of model type.
@@
 For a new TQA instance, we initialize the memory state $s_{0}=(\text{table},\text{question},\text{texts})$, i.e., with the input table, the question, and (if given) textual context.
-[^1] Some TQA datasets, such as TAT, provide textual context in addition to tables and questions.
 All parts of the memory are represented as strings.
@@
 Intent Instruction: Format and Example Tool Selection Code Generation Tool Use/Execution
-Retrieval textual description of what to retrieve, e.g., “sell numbers of 2019” $t=Python$ $M_{c}$ Python interpreter is run on generated code
+Retrieval textual description of what to retrieve, e.g., “sale numbers of 2019” $t=Python$ $M_{c}$ Python interpreter is run on generated code
 Calculation formula or textual description of what to calculate, e.g., “(135-114)/135” $t=T_{cal}$ if formula, else $t=T_{python}$ $M_{c}$ if not formula Calculator is executed or Python interpreter is run on code
@@
 Instructions that require calculation, counting or comparison are captured by Calculation.
-To fulfill the possible need for external (factual) knowledge that is not present in the table or textual context, we add the intent Search.
+To fulfill the possible need for external (factual) knowledge that is not present in the table or textual context, we add the intent Search, which performs Wikipedia searches to retrieve informative text passages.
 Read covers the need for contextual reasoning in table-text QA.
@@
 The intent Finish stops $M_{p}$ from generating more actions and ends the iterative execution of our framework, providing the final answer in the corresponding instruction.
-Lastly, we use an intent called Ask to retrieve an answer based on the internal knowledge from the planning agent.
+Lastly, we use an intent called Ask to retrieve an answer based on the internal knowledge of the planning agent.
+If $M_{p}$ fails to generate a valid action at an iteration, it will continue to generate the action at the next iteration until reaching the maximum iteration number $I$, and return the most common prediction directly from $M_{p}$ as the final answer (See section 3.7).
 3.2 Action Selection
@@
 Note that the instruction of Calculation can also be a textual description, such as “Compute the average number of medals for each country in the table.” To better address these instructions, we introduce a coding agent $M_{c}$ with a Python interpreter, denoted as $T_{python}$ (see 3.4).
+We do not distinguish formulas and text description.
+This means any instructions with the intent Calculation will be firstly passed to $T_{cal}$.
+If $T_{cal}$ fails to execute the instruction, $T_{python}$ is applied.
 The intent Retrieval is also addressed by $M_{c}$ and $T_{python}$, i.e., $M_{c}$ generates Python code based on a given instruction to retrieve target cells in the tables and the Python interpreter returns the executed results.
@@
 We categorize the recent TQA systems into two groups indicating if they require LLM fine-tuning or not.
-As systems requiring fine-tuning, we include OmniTab (Jiang et al., 2022), TableLlama Zhang et al.
+We include the following baselines that require fine-tuning: OmniTab (Jiang et al., 2022), TableLlama Zhang et al.
 (2024a), ProTrix (Wu and Feng, 2024), TAT-LLM (Zhu et al., 2024) and TableLLM Zhang et al.
@@
 Except for OmniTab, which is backboned by BART (Lewis et al., 2020), all others build on top of LLaMA 7b (Touvron et al., 2023).
-As TQA systems that do not require fine-tuning, we include Dater (Ye et al., 2023), Binder (Cheng et al., 2023), Chain-of-Table (Wang et al., 2024), ReAcTable (Zhang et al., 2024c), TabSQLify (Nahid and Rafiei, 2024), Plan-then-Reason (Wu and Feng, 2024), Mix-SC (Liu et al., 2023) and ARC Zhang et al.
+Our set of TQA baselines that do not require fine-tuning includes Dater (Ye et al., 2023), Binder (Cheng et al., 2023), Chain-of-Table (Wang et al., 2024), ReAcTable (Zhang et al., 2024c), TabSQLify (Nahid and Rafiei, 2024), Plan-then-Reason (Wu and Feng, 2024), Mix-SC (Liu et al., 2023) and ARC Zhang et al.
 (2023b).
-They all use GPT-3.5-turbo.
+They all rely on GPT-3.5-turbo.
 Experimental settings.
 In MACT, the choice of the planning and coding agent is flexible as no fine-tuning is involved.
-We experiment with the best open-weight LLMs available at the time of writing: Qwen 72B (Yang et al., 2024) (planning agent) and CodeLLaMA-34B (coding agent).
+We experiment with the best open-weight LLMs available at the time of writing: Qwen-2 72B (Yang et al., 2024) (planning agent) and CodeLLaMA-34B (coding agent).
 They are run on int4 and full precision, respectively.
@@
 Second, we examine how our method performs compared to fine-tuned open-weight LLMs.
+[^3] We did not run ARC as no code is available.
+Results for TAT-LLM is only reported on TAT as the model is specially designed for TAT dataset that features both tables and texts as inputs.
 Following prior work, we use exact match (EM) as the evaluation measure.
@@
 Humans (Crowdsourcing) - 84.1 - 84.7
-Table 3: Results of models without fine-tuning.
+Table 3: Exact Match results of models without fine-tuning.
 SCT refers to SCITAB.
-The models are grouped by the LLM they use as their backbone (top: GPT-3.5; bottom: open-weight LLMs as indicated in parentheses).
+The models are grouped by the LLM they use as their backbone (top: GPT-3.5; middle: open-weight LLMs as indicated in parentheses).
 Performances marked with * are taken from the original paper.
@@
 (2023) for each dataset.
+We bold the best performances in each group.
 MACT outperforms out-of-the-box open-weight LLMs across datasets, demonstrating the effectiveness of specialized agents.
-The middle part of Table 3 provides the results of MACT using specialized agents (MACT (Qw+CL): Qwen as planning agent, CodeLLaMA as coding agent).
+The middle part of Table 3 provides the results of MACT using specialized agents (MACT (Qw+CL): Qwen-2 as planning agent, CodeLLaMA as coding agent).
 As baselines, we compare a setting without a specialized coding agent (MACT (Qw+Qw)) and a setting without a general planning agent (MACT(CL+CL)).
@@
 MACT (Qw+CL)) improves performance considerably.
+MACT with open-weight models delivers comparable performance as closed-source systems.
+As shown in Table 3, by comparing MACT (Qw+CL) with TQA systems that rely on closed-source LLMs (in the upper part of Table 3), we find that our model outperforms the examined TQA systems for three out of four datasets.
+Our multi-agent TQA system is more cost-efficient and straightforward to replicate, while delivering superior performance compared to closed-source TQA models.
+In addition, we show two upper-bounds in the bottom part of Table 3: The performance of human annotators and directly prompting GPT-4 with the table and question.
+As shown in Table 3, GPT-4 has an advantage on TAT and SCITAB.
+On WTQ, we observe comparable performances between MACT (QW+CL) and GPT-4.
+On CRT, our method even outperforms GPT-4 by 5.7%.
+CRT is the most complex dataset, requiring multi-step and multi-category reasoning, which direct inference with GPT-4 cannot generally solve.
+Our step-wise collaborative planning setting is well-suited to such settings.
+In contrast, there is a large gap between MACT and human performance in SCITAB.
+SCITAB collects data from scientific papers, in which abbreviations and domain-specific terms are common.
+These can pose challenges to current systems and models.
+In TAT, MACT often finds the correct answer but struggles to output it in the correct format (see Sec.
+6).
+MACT generalizes better across datasets than fine-tuned TQA systems.
+Table 4 compares our framework with prior fine-tuned TQA models.
+We present results for MACT with different planning and coding agents: Our standard setting (LlaMA-7b+CS-7b) as well as with a stronger planner (Qw-7b+CS-7b).
+In general, for fine-tuned TQA models, their performance on the dataset used for fine-tuning is rather high while they suffer from a considerable drop in EM when tested on other datasets.
+This observation is in line with the findings by Zhang et al.
+(2024a) and Huang et al.
+(2024).
+In contrast, MACT does not use fine-tuned models and can, thus, be applied to any dataset with a good generalization performance.
+MACT demonstrates comparable results to Protrix when using LlaMA-7b as the planning agent, though it has not been fine-tuned.
+As expected, using a better planning agent leads to better results.
+This also shows the robustness of MACT in terms of backbone models.
 c WTQ TAT CRT SCT
 OmniTab (BART 406m) c 62.3* 17.1 20.6 29.1
-TableLlama (LlaMA 7b) c 29.9 17.4 26.9 38.6
-Protrix (LlaMA 7b) c 48.9* 26.8 40.2 42.4
-TAT-LLM (LlaMA 7b) c - 69.6* - -
+TableLlama (LlaMA-7b) c 29.9 17.4 26.9 38.6
+Protrix (LlaMA-7b) c 48.9* 26.8 40.2 42.4
+TAT-LLM (LlaMA-7b) c - 69.6* - -
 MACT (LlaMA-7b+CS-7b) c 38.1 28.3 40.0 41.1
 MACT (Qw-7b+CS-7b) c 58.4 61.9 46.4 45.9
-Table 4: Results of MACT using different LLM agents in comparison to fine-tuned TQA models.
+Table 4: Exact Match Results of MACT using different LLM agents in comparison to fine-tuned TQA models.
 Performances marked with * refer to the in-domain setting (where fine-tuning took place).
@@
 CS refers to the deepseek-coder model.
-MACT with open-source models delivers comparable performances as closed-source systems.
-As shown in Table 3, by comparing MACT (Qw+CL) with TQA systems that rely on closed-source LLMs (in the upper part of Table 3), we find that our model outperforms the examined TQA systems for three out of four datasets.
-This is an important finding, as TQA systems based on commercial closed-source LLMs are costly to operate and often lack reproducibility.
-In contrast, our multi-agent TQA system is more cost-efficient and straightforward to replicate, while delivering superior performance compared to closed-source systems.
-In addition, we show two upper-bounds in the bottom part of Table 3: The performance of human annotators and of GPT-4.
-We observe that MACT achieves comparable performance as GPT-4 on two datasets, without the need for fine-tuning.
-MACT generalizes better across datasets than fine-tuned TQA systems.
-Table 4 compares our framework with prior fine-tuned TQA models.
-We present results for MACT with different planning and coding agents: Our standard setting (LlaMA-7b+CS-7b) as well as with a stronger planner (Qw-7b+CS-7b).
-For the models from related work, we mark the dataset on which they have been fine-tuned with * signs.
-In general, their performance on the dataset used for fine-tuning is rather high while they suffer from a considerable drop in EM when tested on other datasets.
-This observation is in line with the findings by (Zhang et al., 2024a; Huang et al., 2024).
-In contrast, MACT does not use fine-tuned models and can, thus, be applied to any dataset with a good generalization performance.
-MACT demonstrates comparable results as Protrix when using LlaMA-7b as the planning agent, though it has not been fine-tuned.
-As expected, using a better planning agent leads to better results.
-This also shows the robustness of MACT in terms of backbone models.
-MACT is moderately efficient.
-Table 5 compares MACT with other approaches in terms of the total number of prompting (generation) for each instance.
+MACT adapts computational cost to instance complexity.
+Table 5 compares MACT with other approaches in terms of the total number of LLM calls for each instance.
 For Binder and Dater, SC is performed a fixed number of times regardless of problem complexity.
-This results in a high number of prompting per instance, thus being inefficient.
+This results in a high number of LLM calls per instance, making them inefficient.
 In contrast, MACT provides flexibility in generation, as the number of iterations depends on the problem’s complexity.
-For instance, most questions can be solved within three steps for WTQ (see our analysis in A.3), this results in a total of 25 times prompting3332 steps involve action and execution generation, with each five times, plus last step five times of action generation: 2*(5+5)+5.
-at most for each instance.
-If we incorporate the efficiency optimization module, which potentially saves up to one-third of the iterations (see Section 6), the total number of generations per instance is even lower (approximately 15), making MACT comparable to other approaches in terms of efficiency.
-The iterative nature of MACT can lead to a higher upper-bound of generation.
-However, it also allows for more complex problems, making the approach more tailored to real-life needs.
-c Prompting number per instance
+For instance, most questions can be solved within three steps for WTQ (see our analysis in A.3).
+This results in a total of at most 25 LLM calls 4442 steps involve action and execution generation, with each five times, plus last step five times of action generation: 2*(5+5)+5.
+for each instance.
+If we incorporate the efficiency optimization module, which potentially saves up to one-third of the iterations (see Section 6), the total number of LLM calls per instance is even lower (approximately 15), making MACT comparable to other approaches in terms of efficiency.
+The iterative nature of MACT can lead to a higher upper-bound of LLM calls.
+However, it also allows for solving more complex problems, making the approach more tailored to real-life requirements.
+c Number of LLM calls per instance
 Binder c 50
@@
 MACT c 5-65
-Table 5: Prompting number of different approaches. We show lower and upper-bound if not deterministic.
-6 Analysis
-We conduct various analyses of our framework to back up our claims and contributions.
-Unless mentioned otherwise, all analyses are performed using Qwen-72b as $M_{p}$, CodeLlama-34b as $M_{c}$, the number of action generation $k=5$, and selection model $f_{s}=\texttt{SC}$.
-To explicitly analyze the effects of multi-agent collaboration with tool use, we do not use efficiency optimization, which means all instances undergo the iterative collaboration between $M_{p}$ and $M_{c}$ with tool use.
-Further analysis to support our choices of the sampling size $k$ and the maximum number of iterations $I$ are in A.2 and A.3.
+Table 5: Number of LLM calls for different approaches. We show lower and upper-bounds if not deterministic.
 WTQ TAT CRT SCITAB
@@
 Table 6: Ablation study. $T_{srh}$ and $T_{cal}$ refer to the Wikipedia search tool and the calculator tool. $T$ includes the above two tools and a Python interpreter. $M_{c}$ is the coding agent.
+6 Analysis
+We conduct various analyses of our framework to back up our claims and contributions.
+Unless mentioned otherwise, all analyses are performed using Qwen-2 72B as $M_{p}$, CodeLlama-34B as $M_{c}$, the number of action generation $k=5$, and selection model $f_{s}=\texttt{SC}$.
+To explicitly analyze the effects of multi-agent collaboration with tool use, we do not use efficiency optimization, which means all instances undergo the iterative collaboration between $M_{p}$ and $M_{c}$ with tool use.
+Further analysis to support our choices of the sampling size $k$ and the maximum number of iterations $I$ are in A.2 and A.3.
+A case study can be found in 5.
+Figure 2: Distribution of action intents by dataset.
 Effect of Multi-Agent-Collaboration with Tool Use.
@@
 For instance, ablating the search tool barely influences the results whereas there are large performance drops when further ablating the coding agent and the Python interpreter.
-This might be due to the frequency of the usage of tools and the coding agent.
 We find that the search tool is barely used whereas the coding agent is called in almost every query.
-Detailed analysis can be found in A.5.
-In addition, since Wikipedia is a common pre-trained corpus for LLMs, most information might have already been encoded.
+Since Wikipedia is a common pre-training corpus for LLMs, most information might have already been encoded in the LLM.
 Nevertheless, the search tool can still be helpful given LLMs are known to suffer from hallucinations and the knowledge encoded might not be updated in time.
+For more specialized domains and sources, the search tool may be crucial.
 We further observe that the ablation affects WTQ and TAT more than CRT and SCITAB.
@@
 By evaluating our framework on instances from CRT that have answers other than yes/no, we find a performance drop of 8.23 when ablating both tools and the coding agent.
+Figure 3: EM (line chart) and iteration ratio (bar chart) against different $\alpha$.
+The iteration ratio is calculated by dividing the number of iterations when using efficiency optimization with a specific $\alpha$ by the number of iterations when not using it (without shortcuts).
+Analysis of Intent Distribution.
+We report the distribution of the intents of the selected actions for each dataset in Figure 2.
+We observe that Retrieve and Calculate are the most frequent intents, along with Finish.
+This indicates that our proposed second agent, the coding agent $M_{c}$ is used frequently.
+Different datasets also require different intents.
+In particular, the framework needs to use the intent Read for solving instances in TAT, where textual descriptions are given while this is not the case for the other datasets.
+We notice that the Search intent is used only few times across datasets.
+This might be because most instances were designed to be solved using the given table and text information.
+However, when looking into the individual cases where Search is used, we still find it useful.
+For instance, one question from WTQ asks about the number of athletes from American but no information about nationality is given in the table.
+In this case, Search assists with answering the question by adding the nationality for each athlete from Wikipedia.
+Though the intents Read, Search and Ask are less used compared to others, we still incorporate them to adapt to various use cases that might occur in real-life use cases.
 Effect of Efficiency Optimization.
-Figure 2: EM (bar chart) and iteration ratio (line chart) against different $\alpha$.
-The iteration ratio is calculated by dividing the number of iterations when using efficiency optimization with a specific $\alpha$ by the number of iterations when not using it.
-To investigate the trade-off between efficiency and accuracy, we plot the model performance against $\alpha\in\{0.2,0.4,0.6,0.8,1\}$.
+To investigate the trade-off between efficiency and accuracy, we plot the model performance against $\alpha\in\{0.2,0.4,0.6,0.8,1\}$ in Figure 3.
 We also plot the ratio of the total number of iterations taken to terminate from each tested $\alpha$ to the total number of iterations taken when not using the optimization, i.e., letting the planning agent decide via the Finish action when to stop the execution.
-Figure 2 shows that the best performance is reached for $\alpha=1$, i.e., when requiring all estimated results to agree with each other in order to stop the iteration.
-For SCITAB, for instance, we save approximately 40% of the iterations when setting $\alpha$ to 1 without losing performance compared to not using the optimization component (59.8 vs.
-59.7) On average, adding the efficiency optimization module saves up to 33% of iterations.
-This shows the effectiveness of the optimization in general and that users can individually tune the desired trade-off of performance and computation time.
-Error Analysis
+The best performance is reached for $\alpha=1$, i.e., when requiring all estimated results to agree with each other to stop the iteration.
+For SCITAB, for instance, we save approximately 40% of the iterations when setting $\alpha$ to 1 without losing performance compared to not using the optimization component (59.8% vs.
+59.7%).
+On average, adding the efficiency optimization module saves up to 33% of iterations.
+This shows the effectiveness of the optimization and that users can individually tune the desired trade-off of performance and computation time.
+Error Analysis.
 We randomly sample 50 instances that MACT fails per dataset and conduct an error analysis.
-About half of the errors come from the invalid or wrong code generated by the coding agent $M_{c}$.
-Either $M_{c}$ fails to understand instructions, thus generating wrong code, or due to the complex table data type, the code execution is not successful.
-This indicates the importance of table pre-processing.
+About half of the errors come from invalid or wrong code generated by the coding agent $M_{c}$.
+Either $M_{c}$ fails to make sense of instructions or of complex table structure.
 The second error type can be attributed to evaluation.
-We find that about one-third of failures come from strict evaluation metrics (Exact Match Accuracy).
-This influences the performance of MACT on the TAT dataset the most, as it features long text strings as answers.
+We find that about one-third of failures come from strict evaluation metrics (EM).
+This influences the performance of MACT particularly on the TAT dataset, as it features long text strings as answers.
 The evaluation challenge has been discussed in many previous works Wu and Feng (2024); Li et al.
 (2024).
-To estimate the upper-bound of our method, we use GPT-4 for evaluation and achieve an accuracy of 87.8% on the TAT dataset (66.2% with EM).
-The remaining error cases can be largely attributed to the failure of the planning agent.
-This means the planning agent does not decompose questions correctly.
-For instance, one question asks the score range (min-max) for the top 10 finishers.
+To estimate the upper-bound of our method, we use GPT-4 as evaluator to determine if a predicted answer is semantically the same as the reference answer.
+This results in an accuracy of 87.8% on the TAT dataset, compared to 66.2% with EM.
+The remaining error cases can be largely attributed to the failure of the planning agent in decomposing questions correctly.
+For instance, one question asks for the score range (min-max) of the top 10 finishers.
 Apart from retrieving the min and max scores of the top 10 finishers, the planner continues to generate the action: Calculate[Calculate the range of the scores in the observation 1.].
-This results in the wrong final prediction of the difference between the min and max scores.
+This leads to a wrong prediction.
 7 Conclusions
-We propose multi-agent collaboration with tool use for table question answering.
-Unlike previous work, our MACT framework neither requires fine-tuning nor does it depend on closed-source models.
-In our experiments, our framework demonstrates generalizability across different benchmark datasets and outperforms a number of state-of-the-art approaches, including closed-source commercial models and fine-tuned models.
+We have proposed MACT, a multi-agent collaboration with tool use for table question answering.
+Unlike previous work, MACT neither requires fine-tuning nor does it depend on closed-source models.
+In our experiments, our framework demonstrates good generalizability across different benchmark datasets and outperforms a number of state-of-the-art approaches, including closed-source commercial models and fine-tuned models.
 To boost efficiency, we introduce an efficiency optimization module that saves up to 33% of the iterations in our analysis.
 In our experiments and analyses, we show that multi-agent collaboration with tools is an effective approach for table question answering.
-8 Limitation
+8 Limitations
 MACT is evaluated mainly with single table settings due to the scarcity of datasets featuring multi-table complex reasoning.
@@
 Secondly, we only study TQA in the context of English, while there exist many multi-lingual TQA benchmarks and challenges.
+Acknowledgements
+This work was partially supported by the EU Project SMARTY (GA 101140087).
 Appendix A Appendix
@@
 (2023b) and SCITAB (Lu et al., 2023) are publicly available under the licenses of CC-BY-SA-4.05, MIT and MIT, respectively.
-These licenses all permit us to compose, modify, publish, and dis- tribute additional annotations upon the original dataset.
+These licenses all permit us to compose, modify, publish, and distribute additional annotations upon the original dataset.
 Dataset #Test M-step M-category Domain
@@
 Overall, this analysis confirms that SC as selection model is a good choice.
-A.5 Analysis of Intent Distribution
-We report the distribution of the intents of the selected actions for each dataset in Figure 3.
-We observe that Retrieve and Calculate are the most frequent intents, along with Finish.
-This indicates that our proposed second agent, the coding agent $M_{c}$ is used frequently.
-Different datasets also require different intents.
-In particular, the framework needs to use the intent Read for solving instances in TAT, where textual descriptions are given while this is not the case for the other datasets.
-We notice that the Search intent is used only few times across datasets.
-This might be because most instances were designed to be solved using the given table and text information.
-However, when looking into the individual cases where Search is used, we still find it useful.
-For instance, one question from WTQ asks about the number of athletes from American but no information about nationality is given in the table.
-In this case, Search assists with answering the question by adding the nationality for each athlete from Wikipedia.
-Figure 3: The distribution of action intents for different datasets.
+A.5 Case Study
+We present two reasoning traces selected from CRT and WTQ in Figure 4 and 5, respectively.
+Figure 4 presents a case involving complex reasoning, where models need to identify top ten finishers, calculate percentage of drivers for each constructor and return the constructor with the highest percentage.
+In Figure 5, we observe the planning agent has the ability to self-correct in Thought 3.
+Figure 4: An instance selected from CRT featuring complex reasoning.
+Figure 5: An instance selected from WTQ. We find the planning agent can perform self-correct given previous reasoning traces.
 A.6 MACT Prompts
-We provide the prompts used for the planning agent for the examined datasets (WTQ, TAT, CRT and SCITAB) in Figures 6, 7, 8, and 9.
+We provide the prompts used for the planning agent for the examined datasets (WTQ, TAT, CRT and SCITAB) in Figures 8, 9, 10, and 11.
 Figure 12 and Figure 13 show the prompts used for the coding agent for the action intents Retrieval and Calculation, respectively.